OpenStudy (anonymous):

what is the average rate of change of the function f(x) - 4x-7 between x=-1 and x = 2

OpenStudy (anonymous):

(4-7-(-8-7))/3=-12/3=-4

OpenStudy (anonymous):

how did you do this

OpenStudy (anonymous):

rate of change for a line is generally called the "slope" and it is \(-4\) in this case since your line looks like \(y=mx+b\) with \(m=-4\)

OpenStudy (anonymous):

well, in order to solve this, you have to use the def:( f(b)-f(a))/(b-a) to get the right answer